OpenStudy (anonymous):

Grace is making party hats for her upcoming birthday party. A single hat has a diameter of 21 centimeters and a slant height of 38 centimeters. What is the lateral area of each party hat? Answer 3,892.43 cm2 2,506.99 cm2 1,599.86 cm2 1,253.50 cm2

sam (.sam.):

\[(\pi)(radius)(slant height) = (\pi)(10.5)(38) = 399(\pi)cm^2=1,253.50 cm ^{2}\]

OpenStudy (anonymous):

ok so do you not have to multiply it by 2?? i dont understand.. i thought that 2pi rh ?

sam (.sam.):

That's a formula you have to remember.